# Vulnerabilities that arise from poor governance in Distributed Ledger Technologies

### v1

@@ 4. Governance Properties
4.0.1. Participation
First we begin by considering who may participate in the decision making process of DLTs. These actors can be classified into the following set of agents: users, service providers, proposers and decision-makers.
@@ 4.0.1. Participation
We advocate for equal suffrage. This property removes the possibility of weighted votes, thereby enforcing equal voting power distribution.
@@
Possible implementations: Property P1 advocates for a one-person-one-vote mechanism to distribute voting power. This can be easily achieved in permissioned DLTs, since each participant’s identity is known. In permissionless DLTs, this is more challenging to achieve if a decentralised structure is to be retained [^1] It can be argued that some permissionless DLTs are already not truly decentralised if the majority of voting power is held by two or three entities and participation rights are assigned proportional to wealth. See Figure 1 for a distribution of hashrate in Bitcoin (regulated through Proof-of-Work) and Figure 2 for the staking pool distribution of Ethereum (regulated through Proof-of-Stake)..
Figure 1. Bitcoin’s hashrate distribution for the year of 2024 to 2024. Source: Blockchain.com
@@
This is because it is non-trivial to distribute unique identity credentials in a decentralised manner, and in a way in which a single user cannot easily obtain multiple identity credentials. Indeed, an ongoing research effort is dedicated to answering this question (Dunphy et al., 2018), (Goodell and Aste, 2019). The Proof of Humanity project aims to address this challenge. It is a peer-to-peer protocol that issues unique identity credentials to its participants(PoH, 2021). They each create an identity, and it is then approved by other validated participants in the protocol. It offers the benefit that it is a more decentralised solution (it distributes trust), but participants must wait for others to vouch for their profile, and it is built on Ethereum so it is only available on their ecosystem. Bridges are protocols that allow for the transfer of data or transactions across different blockchains, and may be leveraged in this case to enable PoH on other compatible blockchains (Lee et al., 2023).
Technologies such as Semaphore, which is an Ethereum based zero-knowledge protocol, allows members to cast a vote without revealing their identity, whilst demonstrating they belong to a group of valid voters (Semaphore projects, 2022). In this manner, voters can only participate if they prove in zero-knowledge that they own a valid, unique voter credential, and can vote without having to reveal it. It is possible to assign one person one vote in permissionless DLTs by combining a protocol such as Proof of Humanity with Semaphore. Proof of Humanity ensures only approved (real) individuals receive a unique identity credential. Then, voters can prove they own such a credential in zero-knowledge, meaning they do not need to publicly disclose their credential, through Semaphore. The work in (Gangemi and Kharman, 2023) proposes combining these two protocols in the context of resolving disputes is Web 3 platforms. We note both Proof of Humanity and Semaphore are only compatible with Ethereum. We encourage DLT engineers and researchers to prioritise delivering these functionalities across varying DLT architectures.
@@
4.0.2. 4.2 - Incentives
Many DLT protocols assume their governance system is secure due to the financial incentives offered to their participants. Unfortunately, in a number of cases these governance protocols are not clearly defined, documentation is sparse, outdated or spread across multiple sources (Foundation, [n. d.]), (IOTA, 2023), (IOTA governance forum, [n. d.]). Some protocols are so complex it is non-trivial to outline a clear mathematical model of how all of the incentives interact and what the net best strategy for an agent would be (Polkadot, 2023). As a result, undesirable scenarios arise whereby agents are inadvertently incentivised to undermine the network (Zook and Blankenship, 2018). There are two crucial incentives that must be clearly defined and identifiable in the governance protocol:
@@
4.0.3. 4.3 - Voting Mechanism
@@
This does not bode well for ordinal voting schemes, but none of these results apply to cardinal voting schemes. A caveat with cardinal voting schemes is that it is important to find a meaningful function to allocate value to preferences. One example is Quadratic Voting (Lalley et al., 2016), whereby each extra vote for a candidate is quadratically more expensive for a voter. Another class of voting mechanisms are those that are probabilistic. An example is the serial random dictator model or Maximum Entropy Voting (Sewell et al., 2009). Probabilistic voting mechanisms offer desirable properties such as V1, V2 and V4 (or slightly weaker notions of the latter two), and in some instances even V3 (Sewell et al., 2009), (Bade, 2015), (Kosheleva et al., 2021), (Bade, 2020). We remind DLT designers of the classical desirable properties of voting schemes, although these can be side-stepped for slightly weaker notions in order to achieve a better compromise. Nonetheless, we strongly advocate for cardinal (provided a meaningful utility function to allocate preference intensity exists) or probabilistic schemes.
4.0.4. 4.4 - Governance Liveness
In the context of blockchain, liveness is the property that ensures nodes in the network are available to exchange information and eventually achieve consensus (Alpern and Schneider, 1985). In the context of governance we define the property of liveness as follows:
  • L1
    liveness in governance: the governance process eventually terminates and returns a valid outcome.
This property is crucial to ensure the system cannot be attacked by refusing to output an outcome. This is especially important in states of exception, for example in the case where a major vulnerability has been discovered and must be patched.
4.0.5. 4.5 - Security
@@
  • S4
    Censorship resistance: it should not be possible to suppress a valid voter’s ballot.
@@
Possible implementations: A number of voting scheme implementations exist that satisfy notions of ballot secrecy. Well studied proposals include Helios(Adida, 2008), JCJ (Juels et al., 2005), Civtas (Clarkson et al., 2008) and Athena (Smyth, 2019). All of these protocols offer encrypted ballots, but not all provably satisfy the same notions of ballot secrecy, some have weaker security (earlier versions of Helios), others stronger, (JCJ and Civtas) but pay a price for it with an increased computational complexity. We note that none of these schemes have been implemented on-chain, although there are numerous voting smart contracts available. Whilst other on-chain voting schemes exist, most do not provide encrypted ballots. We find VoteCoin (Scholtz, 2022) and MACI (Buterin et al., 2023) are the only on-chain voting schemes that provide encrypted ballots. Unfortunately, VoteCoin provides encrypted ballots only while the election is occurring. This forgoes secrecy once the election is over, enabling coercers to check if their victim complied. The same can be said about the work in (Košt’ál et al., 2019). In MACI, votes remain encrypted, but the coordinator tasked with tallying individually decrypts ballots. If these are intercepted or if the coordinator is not trustworthy, votes may be easily revealed. Note that other protocols exist that provide anonymous voting, whereby the vote is public but the identity of the voter is hidden. Examples include Snapshot (Labs, 2023) and Vocdoni (Vocdoni, 2021). This is not congruent to secret ballots and should not be considered an appropriate security measure especially in DLT implementations, where identities are at best pseudonymous.
In summary, DLT practitioners have a number of options to ensure their governance protocol satisfies notions of ballot secrecy. They may run off-the-shelf protocols such as Helios’16, Civtas, Athena or others previously mentioned. These can be run off-chain, and the results may be notarised on-chain leveraging zero-knowledge proofs. These can be used to demonstrate the election result was correctly computed. Alternatively, DLT practitioners may implement these voting schemes on-chain, although this option will incur greater time and cost. We discourage from following this path. The aforementioned voting schemes have been well studied and audited. Implementing these on-chain would require security audits and verifying that no bugs arise or inaccuracies appeared. Alternatively, if on-chain voting is essential, practitioners may settle for options such as MACI, VoteCoin or Vocdoni which offer encrypted ballots, but these options are insecure and neither meet formal notions of ballot secrecy.
@@
  • Individual verifiability: voters can check that their own ballots are recorded.
@@
Possible implementations: regarding available implementations, all e-voting schemes satisfy notions of verifiability (some weaker than others). It becomes a matter of deciding the level of security and the trust assumptions one can make in their system, to pick the right verifiability formalisation. Achieving verifiability alone is trivial, as can be seen from most blockchain voting implementations: simply publicly revealing votes yields verifiable voting schemes. The challenge lies in achieving both verifiability and secrecy, without sacrificing one in favour of the other. We stress: practitioners do not need to compromise, achieving both is indeed possible, and is in fact necessary. The protocols outlined in 4.0.5 provide verifiability as well as ballot secrecy (granted that each adhere to different formal notions of both ballot secrecy and verifiability).
Coercion Resistance: The work in (Juels et al., 2005) introduced a paradigm shift in the concept of coercion-resistance. Prior to it, Benaloh and Tunistra introduce the notion of receipt-freeness as means to ensure voters cannot prove to a thrid party how they voted (Benaloh and Tuinstra, 1994). However, the authors in (Juels et al., 2005) consider a much more powerful adversary, one that may ‘demand of coerced voters that they vote in a particular manner, abstain from voting, or even disclose their secret keys’. They provide a formal notion of coercion resistance, and an election protocol that provably satisfies said notion. Note that this definition of coercion resistance partly captures censorship resistance, as the adversary may demand of the voter to abstain. However, it does not consider an adversary that may suppress a ballot from the tallying phase. A recent survey of coercion resistant voting schemes can be found in (Battagliola et al., 2024) in section 6.5.1, and their respective time complexities is shown in Table 6.2. None of these schemes are implemented on-chain. The authors in (Spadafora et al., 2020) propose a coercion resistant voting protocol using blockchain to cast votes. They also propose a security notion of Vote Indistinguishability (VI), which is congruent to secret ballots, and show their voting protocol is VI secure.
@@
5. Missing properties

### v2

@@ 7. Governance Properties
7.1. Participation
First we begin by considering who may participate in the decision making process of DLTs. These actors can be classified into the following set of agents: users, service providers, proposers and decision-makers. We introduce some terminology for ease of exposition.
@@ 7.1. Participation
We advocate for equal suffrage. This property means everyone eligible to vote has the right to cast a vote, and that everyone’s vote carries the same weight (Shaw, 1914). This property removes the possibility of weighted votes, thereby enforcing equal voting power distribution.
@@
Possible implementations: Property P1 advocates for a one-person-one-vote mechanism to distribute voting power. This can be easily achieved in permissioned DLTs, since each participant’s identity is known. In permissionless DLTs, this is more challenging to achieve if a decentralised structure is to be retained [^11] It can be argued that some permissionless DLTs are already not truly decentralised if the majority of voting power is held by a small number of entities and participation rights are assigned proportional to wealth. See Figure 2 for a distribution of hashrate in Bitcoin (regulated through Proof-of-Work) and Figure 3 for the staking pool distribution of Ethereum (regulated through Proof-of-Stake)..
Figure 2. Bitcoin’s hashrate distribution for the year of 2024. Source: Blockchain.com
@@
This is because it is non-trivial to distribute unique identity credentials in a decentralised manner, and in a way in which a single user cannot easily obtain multiple identity credentials. Indeed, an ongoing research effort is dedicated to answering this question (Dunphy et al., 2018), (Goodell and Aste, 2019). The Proof of Humanity project aims to address this challenge. It is a peer-to-peer protocol that issues unique identity credentials to its participants(PoH, 2021). They each create an identity, and it is then approved by other validated participants in the protocol. It offers the benefit that it is a more decentralised solution (it distributes trust), but participants must wait for others to vouch for their profile, and it is built on Ethereum so it is only available on their ecosystem. Bridges are protocols that allow for the transfer of data or transactions across different blockchains, and may be leveraged in this case to enable Proof of Humanity on other compatible blockchains (Lee et al., 2023).
Semaphore is an Ethereum based zero-knowledge protocol that allows members to cast a vote without revealing their identity, whilst demonstrating they belong to a group of valid voters (Semaphore projects, 2022). In this manner, voters can only participate if they prove in zero-knowledge that they own a valid, unique voter credential, and can vote without having to reveal it. It is possible to assign one person one vote in permissionless DLTs by combining a protocol such as Proof of Humanity with Semaphore. Proof of Humanity ensures only approved (real) individuals receive a unique identity credential. Then, voters can prove they own such a credential in zero-knowledge, meaning they do not need to publicly disclose their credential, through Semaphore. The work in (Gangemi and Kharman, 2024) proposes combining these two protocols in the context of resolving disputes is Web 3 platforms. We note both Proof of Humanity and Semaphore are only compatible with Ethereum. We encourage DLT engineers and researchers to prioritise delivering these functionalities across varying DLT architectures.
@@
7.2. Incentives
Many DLT protocols assume their governance system is secure due to the financial incentives offered to their participants. Unfortunately, in a number of cases these governance protocols are not clearly defined, documentation is sparse, outdated or spread across multiple sources (Foundation, 2017), (IOTA, 2023), (IOTA governance forum, [n. d.]). Some protocols are so complex it is non-trivial to outline a clear mathematical model of how all of the incentives interact and what the net best strategy for an agent would be (Network, 2016). As a result, undesirable scenarios may arise whereby agents are inadvertently incentivised to undermine the network (Zook and Blankenship, 2018). There are two crucial incentives that must be clearly defined and identifiable in the governance protocol:
@@ 7.2. Incentives
Possible implementations: Some DLTs implement the property of distributed liability (I3) by introducing penalties for non-compliance (also known as slashing (He et al., 2023)), which also creates an incentive to participate honestly in the governance protocol. Implementing properties I1 and I2 is straightforward: direct rewards can be granted to participants upon voting or if the proposals they made receive sufficient support and are correctly implemented. Many cryptocurrencies already provide these rewards (Foundation, 2024; Network, 2016; Foundation, 2018b). To implement I4, elections must be held on-chain, and those with the power to implement the election outcome must be able to participate in the election. The challenge lies in being able to characterise how all these incentives interact with each other, and whether other incentives in the cryptocurrency negatively interfere with them. For example, on-chain voting will introduce fees to cast a vote, and this may directly interfere with a financial incentive offered to voters for participating.
7.3. Voting Mechanism
@@
This does not bode well for ordinal voting schemes, but none of these results apply to cardinal voting schemes. A caveat with cardinal voting schemes is that it is important to find a meaningful function to allocate value to preferences. One example is Quadratic Voting (Lalley et al., 2016), whereby each extra vote for a candidate is quadratically more expensive for a voter. Another class of voting mechanisms are those that are probabilistic [^13] A probabilistic voting scheme returns an outcome depending on a certain probability. The same election run multiple times may not always return the same winner (Burden, 1997).. An example is the serial random dictator model or Maximum Entropy Voting (Sewell et al., 2009). Probabilistic voting mechanisms offer desirable properties such as V1, V2 and V4 (or slightly weaker notions of the latter two), and in some instances even V3 (Sewell et al., 2009), (Bade, 2015), (Kosheleva et al., 2021), (Bade, 2020). They also completely sidestep any of the theorems outlined by Arrow, Gibbard (Gibbard, 1973) and Gibbard–Satterthwaite.
Possible implementations: We strongly advocate for cardinal (provided a meaningful utility function to allocate preference intensity exists) or probabilistic schemes. These have been shown to satisfy the highest number of desirable properties by researchers in the field of social choice.
7.4. Security
@@
Possible implementations: A number of voting scheme implementations exist that satisfy notions of ballot secrecy. Well studied proposals include Helios(Adida, 2008), JCJ (Juels et al., 2005), Civtas (Clarkson et al., 2008) and Athena (Smyth, 2019). All of these protocols offer encrypted ballots, but not all provably satisfy the same notions of ballot secrecy, some have weaker security (earlier versions of Helios), others stronger, (JCJ and Civtas) but pay a price for it with an increased computational complexity. We note that none of these schemes have been implemented on-chain, although there are numerous voting smart contracts available. Whilst other on-chain voting schemes exist, most do not provide encrypted ballots. We find VoteCoin (Scholtz, 2022) and MACI (Buterin et al., 2023) are the only on-chain voting schemes that provide encrypted ballots. Unfortunately, VoteCoin provides encrypted ballots only while the election is occurring. This forgoes secrecy once the election is over, enabling coercers to check if their victim complied. The same can be said about the work in (Košt’ál et al., 2019). In MACI, votes remain encrypted, but the coordinator tasked with tallying individually decrypts ballots. If these are intercepted or if the coordinator is not trustworthy, votes may be easily revealed. Note that other protocols exist that provide anonymous voting, whereby the vote is public but the identity of the voter is hidden. Examples include Snapshot (Labs, 2023) and Vocdoni (Vocdoni, 2021). This is not congruent to secret ballots and should not be considered an appropriate security measure especially in DLT implementations, where identities are at best pseudonymous (Androulaki et al., 2013).
In summary, DLT practitioners have a number of options to ensure their governance protocol satisfies notions of ballot secrecy. They may run off-the-shelf protocols such as Helios’16, Civtas, Athena or others previously mentioned. These can be run off-chain, and the results may be notarised on-chain leveraging zero-knowledge proofs. These can be used to demonstrate the election result was correctly computed. Alternatively, if on-chain voting is essential, practitioners may settle for options such as MACI or VoteCoin, which offer encrypted ballots, but these options are insecure and neither meet formal notions of ballot secrecy.
@@
  • Individual verifiability: voters can check that their own ballots are recorded  [^14] The property of individual verifiability is closely related to censorship resistance, but it is a weaker notion than the latter. Individual verifiability allows a voter to detect if their vote has been suppressed, whereas censorship resistance demands that votes cannot be suppressed in the first place..
@@
Possible implementations: regarding available implementations, all e-voting schemes satisfy notions of verifiability (some weaker than others). It becomes a matter of deciding the level of security and the trust assumptions one can make in their system, to pick the right verifiability formalisation. Achieving verifiability alone is trivial, as can be seen from most blockchain voting implementations: simply publicly revealing votes yields verifiable voting schemes. The challenge lies in achieving both verifiability and secrecy, without sacrificing one in favour of the other. We stress: practitioners do not need to compromise, achieving both is indeed possible, and is in fact necessary. The possible implementations of ballot secrecy outlined earlier in 7.4 provide verifiability as well as ballot secrecy (granted that each adhere to different formal notions of both ballot secrecy and verifiability).
Coercion Resistance: The work in (Juels et al., 2005) introduced a paradigm shift in the concept of coercion-resistance. Prior to it, Benaloh and Tunistra introduce the notion of receipt-freeness as means to ensure voters cannot prove to a thrid party how they voted (Benaloh and Tuinstra, 1994). However, the authors in (Juels et al., 2005) consider a much more powerful adversary, one that may ‘demand of coerced voters that they vote in a particular manner, abstain from voting, or even disclose their secret keys’. They provide a formal notion of coercion resistance, and an election protocol that provably satisfies said notion. Note that this definition of coercion resistance partly captures censorship resistance [^15] The property of censorship resistance means it should not be possible to suppress a valid voter’s ballot (Wahrstätter et al., 2024)., as the adversary may demand of the voter to abstain. However, it does not consider an adversary that may suppress a ballot from the tallying phase. A recent survey of coercion resistant voting schemes can be found in (Battagliola et al., 2024) in section 6.5.1, and their respective time complexities is shown in Table 6.2. None of these schemes are implemented on-chain. The authors in (Spadafora et al., 2020) propose a coercion resistant voting protocol using blockchain to cast votes. They also propose a security notion of Vote Indistinguishability (VI), which is congruent to secret ballots, and show their voting protocol is VI secure.
@@
8. Evaluation
We proceed by selecting a diverse range of cryptocurrencies to evaluate varying types of governance models, DLT architectures and functionalities. Bitcoin and Ethereum serve as foundational examples of blockchain systems with informal or off-chain governance processes (Community, 2011), (Foundation, 2015). Cardano is an interesting case-study, as its participants just recently approved a constitution to formalise its governance process [^16] Cardano Constitutional Convention 2024.. IOTA is an example of a Directed Acyclic Graph DLT, as opposed to a blockchain architecture (Popov, 2018). Solana was selected due to its alternative consensus mechanism, Proof of History, as well as its high transaction per second rates and low transaction fees (Yakovenko, 2018). Avalanche is selected due to its consensus mechanism, the integration of subnetworks [^17] The subnetwork architecture is one in which participants validate different subsets of blockchains in the network (Sekniqi et al., 2020b)., and the inclusion of multiple built-in blockchains (Sekniqi et al., 2020a). Terra Luna was included as a case-study of a cryptocurrency that collapsed partly due to its governance structure, which failed to implement effective safeguards or timely interventions to address vulnerabilities in its algorithmic stable-coin model (Liu et al., 2023b), (Briola et al., 2023).
Property Bitcoin Ethereum IOTA Polkadot Cardano Solana Avalanche Terra Luna
  Equal suffrage No No No No No No No No
Non-purchaseable voting power No No No No No No No No
  Incentive to vote None Social Social Monetary Monetary Monetary Monetary Monetary
Incentive to make beneficial proposals None Social Social Monetary Monetary Monetary Monetary Monetary
Distributed liability Yes (Miners) Yes (Miners, Validators) Yes (Users) Yes (Users, Validators) Yes (Users, Validators) Yes (Validators) Yes (Validators) Yes (Both, Organising Foundation)
Binding election outcome No No Yes Yes Yes Yes Yes Yes
  Independence of irrelevant alternatives No No No No No No No No
Pareto efficient Yes Yes Yes Yes Yes Yes Yes Yes
Non-dictatorial Yes Yes Yes Yes Yes Yes Yes Yes
Strategy-proofness No No No No No No No No
  Ballot secrecy No No No No No No No No
Verifiability Yes Yes Yes Yes Yes Yes Yes Yes
Coercion resistance No No No No No No No No
Censorship resistance Yes Yes Yes Yes Yes Yes Yes Partially
Table 1. Governance properties satisfied in varying cryptocurrencies
For each of the selected cryptocurrencies, we survey their governance documentation and review the findings regarding each governance property below. The sources used for this section are the governance documentation pages of each of the cryptocurrencies: Bitcoin (Community, 2011; Nakamoto, 2008), Ethereum (Foundation, 2015, 2017), IOTA (Foundation, 2018a; IOTA, 2023; IOTA governance forum, [n. d.]), Polkadot (Network, 2016), Cardano (Foundation, 2024), Solana (Foundation, 2018b), Avalanche (Foundation, 2019; Sekniqi et al., 2020a) and Terra Luna (Money, 2019).
8.1. Participation
Regarding suffrage, whilst none of the selected cryptocurrencies satisfy equal suffrage, token-based systems still allow all token holders to participate in governance. In that sense, suffrage is not completely denied to any eligible participant. The inequality arises in voting power, not in the right to vote. All selected cryptocurrencies use some variant of token-based participation except Bitcoin, that grants participation rights proportionally to computational power. None deliver non-purchasable voting power: all selected examples use either Proof of Work, Proof of Stake or a variant of either to achieve consensus. In both of these systems, users may acquire greater power in the network by either purchasing more computational power or investing more wealth in the DLT network.
8.2. Incentives
Incentives to vote are delivered by Polkadot, Cardano, Solana, Avalanche and Terra Luna through direct financial compensation to the voters, whereas the incentive in IOTA is indirect. Stakeholders are motivated by the ecosystem’s success. Ethereum does not offer direct recompense to voters, but applications built on the Ethereum ecosystem often do, such as DAOs. Bitcoin offers no financial reward or alternative incentives for participants to vote in the decision-making processes.
Similarly, Bitcoin offers no rewards or incentives to make beneficial improvement proposals. Bitcoin lacks a formal governance process for proposals. Changes are made via Bitcoin Improvement Proposals (BIPs), which require volunteer effort without direct economic incentives for proposers (Community, 2011). No built-in mechanism rewards making beneficial proposals, so participants rely on altruism or indirect benefits, such as network improvement. Ethereum also lacks direct incentives to make proposals, which are submitted as Ethereum Improvement Proposals (EIPs) (Foundation, 2015). IOTA’s governance includes mechanisms for proposing changes, but there are no explicit financial incentives for proposers (Foundation, 2018a). Participants may be motivated indirectly by the chance to improve the protocol or gain reputation within the ecosystem. The remaining cryptocurrencies all have mechanisms to fund winning proposals, with Terra Luna, Cardano and Polkadot providing said funding from their own treasury (Money, 2019; Foundation, 2024; Network, 2016).
Bitcoin and Ethereum rely on voluntary adoption and off-chain agreement, making their governance outcomes non-binding. The rest of selected cryptocurrencies all implement on-chain voting, thereby making the outcome automatically binding.
Regarding liability, Bitcoin miners are responsible for consensus and validation. In Ethereum, miners were responsible when Ethereum used Proof-of-Work for consensus, and validators [^18] Validators are the service providers in DLTs that use Proof-of-Stake for consensus. are responsible since Ethereum shifted to Proof-of-Stake consensus. Participants indirectly manage the DLT in IOTA, since they can be both users and service providers at the same time. Validators and users share liability for governance and consensus in Polkadot, Cardano, Solana and Avalanche. However, Terra Luna operated differently prior to its collapse: both users and the Organising Foundation (Terraform Labs) assumed liability in the governance process.
8.3. Voting Mechanism
The table below summarises the voting aggregation algorithms used in each cryptocurrency. Most use a token-based variant of plurality voting (also known as first-past-the-post or majority voting). Majority voting is known to satisfy Pareto Efficiency and is non-dictatorial, but it is not independent of irrelevant alternatives and therefore is not strategy proof (Maskin, 2020). The exceptions are Bitcoin and Ethereum, neither of which use on-chain voting. Bitcoin nonetheless satisfies the same properties because it also utilises a majority voting scheme, although the 95% threshold is likely to reduce the influence of irrelevant alternatives in practise, compared to the 51% threshold. We are unable to characterise the properties that Ethereum’s voting scheme satisfies because there are no clear quorum or thresholds defined for adoption.
Cryptocurrency Threshold for Proposal Approval
Bitcoin 95% of miner signalling.
Ethereum ‘Rough consensus’.
Solana 50%+1 of staked SOL (quorum rules may apply).
IOTA 50%+1 of Mana (majority-based).
Cardano Planned: 50%+1 ADA + quorum (higher for major changes).
Polkadot Conviction-weighted majority, often ¿66% for critical votes.
Avalanche 50%+1 AVAX (quorum-dependent).
Terra Luna 40% quorum; 50%+1 approval threshold.
Table 2. Voting thresholds for proposal approval across various cryptocurrencies.
8.4. Security
All selected cryptocurrencies reveal the votes publicly, therefore none satisfy notions of ballot secrecy. Conversely, verifiability is obviously satisfied, since all votes are public, everyone can check the election outcome is correct. The public nature of votes however, makes coercion resistance impossible. Because a voter can prove how they voted, they can be bribed or threatened to vote a certain way. On-chain voting means miners or validators are responsible for processing transactions containing votes. They have the power to exclude transactions from the DLT containing votes, and therefore formal notions of censorship resistance cannot be provably guaranteed. In practise, most miners or validators do include valid transactions and follow the protocol, but censorship resistance is not guaranteed: researchers have found miners may still misbehave by reordering, inserting, or censoring transactions for their own gain (Daian et al., 2020b).
8.5. Missing properties
From the evaluation of the selected cryptocurrencies, we find there are a number of currently disregarded good governance properties. These include: